# Intuitionism with Truth Tables: A Decision Procedure for IPL Based on RNmatrices

### v2

@@ 5 Complexity
Let $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ be some nmatrix. The branching factor of $\mathcal{M}$ is the highest $\kappa$ such that, for every n-ary ~$\tilde{\circ}\in\mathcal{O}$, every formula $\alpha_{1},\ldots,\alpha_{n}$ and every valuation $v\in Val(\mathcal{M})$, |~$|\tilde{\circ}(v(\alpha_{1}),\ldots,v(\alpha_{n}))|\leq\kappa$.
For example, in the case of $\mathcal{M}_{IPL}$, we have that disjunction and conjunction have branching factor $1$, while implication and negation have branching factor $2$. Therefore, in this case, $\kappa=2$.
Remark 5.3.
If we consider the valuations in the nmatrix as branches in trees, we can say that $|\mathcal{V}|^{|atoms(\alpha)|}$ gives the number of trees, while $\kappa^{(|sub(\alpha)|-|atoms(\alpha)|)}$ gives the (maximum) number of branches of each tree.
We can also obtain a lower bound by assuming $\kappa=1$, which is the branching factor of deterministic matrices. This gives us
Definition 5.4.
$lb_{\mathcal{M}}(\alpha)=|\mathcal{V}|^{|atoms(\alpha)|}$ (2)
From (5.1) and (5.4), it follows that
$ub_{\mathcal{M}}(\alpha)=\kappa^{(|sub(\alpha)|-|atoms(\alpha)|)}\times lb_{\mathcal{M}}(\alpha)$ (3)
$p$ $p$
$\mathcal{V}_{1}$ $\mathcal{V}_{1}$
$\vdots$ $\vdots$
$\mathcal{V}_{n}$ $\mathcal{V}_{n}$
Table 10: Base case.
Now we shall prove that these are faithful lower and upper bounds for the number of rows.
Theorem 5.5.
Let $\mathcal{M}$ be a nmatrix and $rows(\alpha)$ the set of every row of the truth table for some formula $\alpha$. Then
$lb_{\mathcal{M}}(\alpha)\leq|rows(\alpha)|\leq ub_{\mathcal{M}}(\alpha)$
Proof.
Let $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ be the nmatrix for the language of $\alpha$, where $\mathcal{V}=\{\mathcal{V}_{1},\ldots,\mathcal{V}_{n}\}$. Induction on the complexity of $\alpha$.
Base case If $\alpha=p$, the truth table is depicted in Table 10, which has $|\mathcal{V}|=n$ rows. Now, $|sub(\alpha)|=|atoms(\alpha)|=1$. For this reason, $lb_{\mathcal{M}}(\alpha)=|\mathcal{V}|^{1}=|\mathcal{V}|$ and $ub(\alpha)=\kappa^{0}\times|\mathcal{V}|^{1}=|\mathcal{V}|$. Therefore, $lb_{\mathcal{M}}(p)=|rows(p)|=ub_{\mathcal{M}}(p)$.
IH If $c(\alpha)<c(\alpha^{\prime})$, then the theorem holds for $\alpha$.
$\alpha=\neg\beta$
Let $rows(\beta)=\{v_{0},\ldots,v_{n}\}$ be the list of rows of the nmatrix for $\beta$. Because $atoms(\beta)=atoms(\neg\beta)$, we conclude that $lb_{\mathcal{M}}(\neg\beta)=lb_{\mathcal{M}}(\beta)$. Now, observe that $|sub(\neg\beta)|=|sub(\beta)|+1$. From this, we get $ub_{\mathcal{M}}(\neg\beta)=\kappa^{|sub(\beta)|-|atoms(\neg\beta)|+1}\times lb_{\mathcal{M}}(\neg\beta)=ub_{\mathcal{M}}(\beta)\times\kappa$. By the definition of branching factor, for every $v_{n}\in rows(\beta)$, |¬~$|\tilde{\neg}(v_{n}(\beta))|\leq\kappa$. Therefore, $|rows(\neg\beta)|\leq|rows(\beta)|\times\kappa$. Now, by (IH), $lb_{\mathcal{M}}(\beta)\leq|rows(\beta)|\leq ub_{\mathcal{M}}(\beta)$. Therefore, $lb_{\mathcal{M}}(\neg\beta)\leq|rows(\neg\beta)|\leq ub_{\mathcal{M}}(\neg\beta)$.
[…]
Let $rows(\beta)=\{v_{0},\ldots,v_{n}\}$ be the nmatrix of $\beta$ and let $rows(\gamma)=\{w_{0},\ldots,w_{n}\}$ be the nmatrix of $\gamma$. By (IH), we get the following
$lb_{\mathcal{M}}(\beta)\leq|rows(\beta)|\leq ub_{\mathcal{M}}(\gamma)$ (4)
$lb_{\mathcal{M}}(\gamma)\leq|rows(\gamma)|\leq ub_{\mathcal{M}}(\gamma)$ (5)
Using 4 and 5 and some laws of algebra, we get
$lb_{\mathcal{M}}(\beta)\times lb_{\mathcal{M}}(\gamma)\leq rows(\beta)\times rows(\gamma)\leq ub_{\mathcal{M}}(\beta)\times ub_{\mathcal{M}}(\gamma)$
Now, to construct the truth table for $\alpha$, we combine $rows(\beta)$ with $rows(\gamma)$. In the worst case, $sub(\beta)\cap sub(\gamma)=\emptyset$. In this case, we append, to each row of $rows(\beta)$, a row of $rows(\gamma)$. Then, each row creates $\kappa-1$ fresh rows. Therefore,
$|rows(\alpha)|\leq|rows(\beta)|\times|rows(\gamma)|\times\kappa$ (6)
Moreover,
$|sub(\alpha)|$ $\leq|sub(\beta)|+|sub(\gamma)|+1$ (7)
$|atoms(\alpha)|$ $\leq|atoms(\beta)|+|atoms(\gamma)|$ (8)
Therefore,
$|\mathcal{V}|^{|atoms(\alpha)|}$ $\leq|\mathcal{V}|^{|atoms(\beta)|+|atoms(\gamma)|}$
$=|\mathcal{V}|^{|atoms(\beta)|}\times|\mathcal{V}|^{|atoms(\gamma)|}$
$=lb_{\mathcal{M}}(\beta)\times lb_{\mathcal{M}}(\gamma)$
and
$\kappa^{(|sub(\alpha)|-|atoms(\alpha)|)}\times|\mathcal{V}|^{|atoms(\alpha)|}$
$\leq\kappa^{(|sub(\beta)|+|sub(\gamma)|+1)-(|atoms(\beta)|+|atoms(\gamma)|)}\times|\mathcal{V}|^{|atoms(\alpha)|}$
$=\kappa^{(|sub(\beta)|-|atoms(\beta)|)+(|sub(\gamma)|-|atoms(\gamma)|)+1}\times\left(|\mathcal{V}|^{|atoms(\beta)|}\times|\mathcal{V}|^{|atoms(\gamma)|}\right)$
$=\left(\kappa^{(|sub(\beta)|-|atoms(\beta)|)}\times|\mathcal{V}|^{|atoms(\beta)|}\right)\times\left(\kappa^{(|sub(\gamma)|-|atoms(\gamma)|)}\times|\mathcal{V}|^{|atoms(\gamma)|}\right)\times\kappa$
$=ub_{\mathcal{M}}(\beta)\times ub_{\mathcal{M}}(\gamma)\times\kappa$
Hence,
$lb_{\mathcal{M}}(\beta\#\gamma)\leq|rows(\beta\#\gamma)|\leq ub_{\mathcal{M}}(\beta\#\gamma)$
5.1 Algorithms
Given a formula $\varphi$, to decide if $\varphi$ is valid we have to proceed in two steps: (1) construct the initial matrix by applying the definitions of the multi-operations, and, then, (2) refine the matrix by removing every spurious row. In the previous section, we have proposed a way to measure the space complexity of (1). Now, we will find the time complexity of (2).
5.1.1 Pseudo-codes
In what follows, given an ordered pair $\langle a;b\rangle$,
$\pi_{1}(\langle a;b\rangle)=a$
and
$\pi_{2}(\langle a;b\rangle)=b$
Moreover, $\mathcal{L}$ denotes the set of every intuitionistic formula and $\mathcal{V}=\{\textbf{F},\textbf{U},\textbf{T}\}$.
Algorithm 1 Compatibility procedure. This algorithm receives two valuations, $v$ and $w$, and check if, for every evaluated formula $\alpha$, if $v(\alpha)=\textbf{T}$ then $w(\alpha)=\textbf{T}$ ($v$ is compatible with $w$). We assume that every valuation list is equally sorted.
procedure isCompatible($v,w:list(\mathcal{L}\times\mathcal{V})$)
     $i\leftarrow 0$
     while $i<length(v)$ do
         $a\leftarrow\pi_{2}(v[i])$triangleright Get the value in $v$ on index $i$.
         if $a=$ T then
              $b\leftarrow\pi_{2}(w[i])$triangleright Get the value in $w$ on the same index.
              if $b=$ T then
                  continue
              else
                  return $false$
              end if
         else
              continue
         end if
         $i\leftarrow i+1$
     end while
     return $true$
end procedure
Algorithm 2 Search algorithm. This algorithm receives a formula $A\in\mathcal{L}$, a (partial) valuation $v:\mathcal{L}\rightarrow\mathcal{V}$ and a matrix. It runs over the entire matrix looking for a valuation $w$ that validates $v$, that is, a $w$ such that $v$ is compatible with and such that $w(A)=\textbf{F}$.
procedure validate($A:\mathcal{L},v:list(\mathcal{L}\times\mathcal{V}),matrix:list(list(\mathcal{L}\times\mathcal{V}))$)
     $i\leftarrow 0$
     while $i<length(matrix)$ do
         $w\leftarrow matrix[i]$triangleright Get a valuation function.
         $j\leftarrow 0$
         while $j<length(w)$ do
              $A^{\prime}\leftarrow\pi_{1}(w[j])$triangleright Get a valuated formula.
              if $A^{\prime}=A$ then
                  $wA\leftarrow\pi_{2}(w[j])$triangleright Get $w(A)$.
                  if $wA=$ F then
                       $compatible\leftarrow isCompatible(v,w)$
                       if $compatible$ then
                           return $true$triangleright In this case, $w$ validates $v$.
                       else
                           continue
                       end if
                  else
                       continue
                  end if
              else
                  continue
              end if
              $j\leftarrow j+1$
         end while
         $i\leftarrow i+1$
     end while
     return $false$triangleright In this case, no $w$ validates $v$.
end procedure
Algorithm 3 Refinement algorithm. In this pseudo-code, we assume that $remove$ is a function that receives a matrix and a row, returning the matrix minus the row.
procedure refineMatrix($matrix:list(list(\mathcal{L}\times\mathcal{V}))$)
     $i\leftarrow 0$
     $newmatrix\leftarrow matrix$
     while $i<length(matrix)$ do
         $v\leftarrow matrix[i]$triangleright Get a valuation function.
         $j\leftarrow 0$
         while $j<length(v)$ do
              $A\leftarrow\pi_{1}(v[j])$triangleright Get a valuated formula.
              $vA\leftarrow\pi_{2}(v[j])$triangleright Get $v(A)$.
              if $vA=$ U then
                  $isValid\leftarrow validate(A,v,newmatrix)$
                  if $isValid$ then
                       continue $\triangleright$ In this case, we keep the matrix unaltered.
                  else
                       $newmatrix\leftarrow remove(v,newmatrix)$
                       break
                  end if
              else
                  continue
              end if
              $j\leftarrow j+1$
         end while
         $i\leftarrow i+1$
     end while
     return $newmatrix$triangleright Return the refined matrix.
end procedure
Algorithm 4 In the best case, the refinement algorithm does not remove any row. However, if it does drag at least one row, we have to rerun it, given that it could have been the case that the removed row was used to validate some other rows. For this reason, in the worst case, the refinement algorithm removes every row and one row at once. This auxiliar procedure should address this possibility.
procedure refineAux($matrix:list(list(\mathcal{L}\times\mathcal{V}))$)
     $i\leftarrow 0$
     $nrows\leftarrow length(matrix)$
     while $i<nrows$ do
         $matrix\leftarrow refineMatrix(matrix)$
         $i\leftarrow i+1$
     end while
     return $matrix$
end procedure
Theorem 5.6.
Given as input a matrix with $n$ rows and $m$ columns, the time complexity of refineAux is $O(n^{3}*m^{3})$.
Proof.
Given a list $v$ with length equal to $m$, in the worst case, the loop in $isCompatible$ runs $m$ times. Therefore, the complexity of this function is $O(m)$. Now, while, in the worst case, the external loop in $validate$ runs $n$ times, where $n$ is the number of rows in the matrix, the internal loop runs $m$ times. Because this functions calls $n*m$ times the function $isCompatible$, then the complexity of $validate$ is $O(n*m)*O(m)=O(n*m^{2})$. In the case of $refineMatrix$, the external loop runs $n$ times and the internal loop runs $m$ times. Therefore, it calls $validate$ $n*m$ times, which gives us $O(n*m)*O(n*m^{2})=O(n^{2}*m^{3})$. Finally, the function $refineAux$ calls $refineMatrix$ $n$ times. Therefore, the complexity of $refineAux$ is $O(n)*O(n^{2}*m^{3})=O(n^{3}*m^{3})$. ∎
5.2 Complexity of $\mathcal{R}(\mathcal{M}_{IPL})$
In the valuation of a formula in $\mathcal{M}_{IPL}$, we consider only $\mathbf{F}$ and $\mathbf{T}$ in the atomic cases. Given that, in this case, $\kappa=2$, we can simplify the upper bound to:
$ub_{\mathcal{M}_{IPL}}(\alpha)$ $=2^{(|sub(\alpha)|-|atoms(\alpha)|)}\times 2^{|atoms(\alpha)|}$ (9)
$=2^{|sub(\alpha)|}$ (10)
Now, given a formula $\varphi\in For(\Omega)$, assume that $|sub(\varphi)|=m$ and that $ub_{\mathcal{M}_{IPL}}(\varphi)=n$. Then
$n^{3}$ $=(2^{m})^{3}=8^{m}$
Therefore, given a formula $\varphi$, the time complexity to compute the truth table (in the worst case) in $\mathcal{R}(\mathcal{M}_{IPL})$ is given by $O(8^{m}*m^{3})$, which is in EXPTIME.
6 Related work and concluding remarks
In this paper, we proposed a new decision procedure for intuitionistic propositional logic IPL based on a 3-valued restricted Nmatrix semantics. Starting from the nice decision procedure for S4 obtained by Grätz in [4] based on a 3-valued RNmatrix, and taking advantage of the conservative translation of Gödel from IPL to S4, it is immediate to obtain an algorithm to decide validity in IPL. By means of an intermediate translation mapping from the language of IPL into the language of S4 we define a 3-valued Nmatrix which is the basis for an RNmatrix semantics for IPL. Moreover, by adapting Grätz’s algorithm for S4, a very simple algorithm for deleting spurious rows in the non-deterministic truth-table generated by the Nmatrix of IPL for a given formula is also provided. The sound and completeness proofs for the RNmatrix and for the obtained truth-tables algorithm for IPL, are presented with full technical details. These proofs were adapted from new ones for S4 we obtained in Subsection 3.2. On the one hand, we reduced the original Nmatrix for S4 presented by Grätz. On the other, we expanded it by adding the truth-tables for conjunction and disjunction. This required the extension and modification of the proofs originally given in [4]. It is worth noting that the Nmatrix for IPL we introduced in Definition 4.10 has a clear and intuitive interpretation in terms of constructibility, by observing the definition of the canonical valuation associated to a $\varphi$-saturated set in IPL given in Definition 4.19, as well as the observations made in Remark 4.20.
The idea of using finite non-deterministic truth-tables with restricted valuations for characterizing intuitionistic logic IPL is not new. Andréa Loparic already obtained in 1977, announced in [7] and finally published in [8], a semantical characterization of Johannson-Kolmogorov minimal logic and intuitionistic propositional logic IPL in terms of a $2$-valued, non truth-functional valuation semantics which induces a decision procedure for these logics. Similar to Grätz method (and to the adaptation to IPL proposed here), Loparic’s approach is based on a criterion for distinguish between that she calls semivaluations and valuations, the latter being the useful ones. This would correspond to level valuations, since they are exactly the semivaluations associate to $\varphi$-saturated sets, and the ones which characterizes the logic. The associated algorithm, as in the cases of S4 and IPL presented in this paper, generates a non-deterministic truth-table for a given formula in which, by using some formal criteria, some rows are deleted, and so the formula is tautology of the given logic if and only if it gets the value 1 in all the undeleted rows.
More recently, Alejandro Solares-Rojas proposed a 3-valued Nmatrix semantics for IPL based on intuitionistic Kripke models and the novel notion of depth-bounded approximations (see [11, Chapter 5]). Within this semantic approach to IPL, a 3-valued Nmatrix – whose truth-values are intuitionistic truth (1), falsity (0) and indeterminacy ($u$) – gives the meaning of the connectives in terms of information that a given agent holds. The valuations over the Nmatrix intend to model actual information states, which are ordered by a certain refinement relation. Such refinement relation aims to model the way knowledge of the agent evolves when new information, provided by reliable external sources, is acquired. The semantics for IPL is given by intuitionistic Kripke frames expanded by a family of valuations indexed by the worlds and where the refinement order corresponds to the accessibility relation of the frame. This interesting 3-valued non-deterministic semantics for IPL departs from a different perspective than ours. Besides the 3-valued Nmatrices proposed in [11, Chapter 5] and the one we give in Definition 4.10 are quite different, it is not immediate whether the semantics proposed by Solares-Rojas could be seen as given by an RNmatrix. This is the subject of future investigation.
It is interesting to notice that, by considering finite and decidable RNmatrix semantics (as the ones for S4 and IPL analyzed here) as a generalization of the standard notion of finite-valued (deterministic) truth-tables, the results presented here for S4 and IPL are a way to overcome the limiting results obtained by Gödel and Dugundji for IPL and S4, respectively, mentioned in Section 1.
As a topic of future research we intend to explore the meaning of the partial valuations corresponding to the rows of the non-deterministic truth-tables in IPL and S4. In particular, we plan to investigate their possible connection with the mini canonical models considered by Hughes and Cresswell in [5, Chapter 8] to prove the finite model property for some modal systems, including S4.
Finally, the extension of this methodology to other logics such as S5 and Prawitz’s ecumenical systems (see [10]) will also be investigated.
Acknowledgments: The first and second author thanks the support of the São Paulo Research Foundation (FAPESP, Brazil) through the PhD scholarship grant #21/01025-3 and the Thematic Project RatioLog #20/16353-3. The second author also acknowledges support from the National Council for Scientific and Technological Development (CNPq, Brazil), through the individual research grant #309830/2023-0.

### v4

@@ 10 Some considerations on complexity
Let $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ be some Nmatrix over a signature $\Theta$. The branching factor of $\mathcal{M}$ is the lowest $\kappa$ such that, for every n-ary $\in\Theta_{n}$ and every $a_{1},\ldots,a_{n}\in\mathcal{V}$, $|\mathcal{O}(\#)(a_{1},\ldots,a_{n})|\leq\kappa$.
Definition 10.2 (Upper bound).
Let $sub(\alpha)$ be the set of all subformulas of $\alpha$, $atoms(\alpha)$ the set of subformulas with complexity equal to 0, $\kappa\geq 1$ the branching factor, and $\theta\leq|\mathcal{V}|$ the number of truth-values used for atomic valuations. Then, the maximum number of rows of the initial matrix for $\alpha$ in $\mathcal{R}(\mathcal{M})$ is given by
$ub_{\mathcal{M}}(\alpha)=\frac{\kappa^{|sub(\alpha)|}}{\kappa^{|atoms(\alpha)|}}\times\theta^{|atoms(\alpha)|}$
Example 10.3.
In $\mathcal{R}(\mathcal{M}_{IPL})$, $\kappa=2$ and $\theta=2$ . Hence,
$ub_{\mathcal{M}_{IPL}}=2^{|sub(\alpha)|}$
A fortiori, we can infer that the upper bound for the Nmatrix is also an upper bound for the restricted Nmatrix. Therefore, we can expect no more than $2^{|sub(\alpha)|}$ rows in a truth-table created for $\alpha$ with $\mathcal{R}(\mathcal{M}_{IPL})$. The truth-table may be shorter than $2^{|sub(\alpha)|}$, but not less than $2^{|atoms(\alpha)|}$, given that there are at least $2^{|atoms(\alpha)|}$ (trivially) valid partial valuations in the truth-table.
This exponential growth in the number of subformulas may be problematic; a formula with only one variable can be arbitrarily large and produce immense tables. However, truth-tables are meaningful. By constructing a truth-table, one generates significant information about a given proposition or argument. A future work direction would be improving the average-case runtime by refining the algorithm with state-of-the-art strategies.
11 Related works
The idea of using finite non-deterministic semantics for characterizing intuitionistic logic and modal logic is not new. This section will mention some approaches to this topic proposed in the literature. We do not intend to present a detailed description and comparison between all these approaches here since they are beyond the scope of the present paper.
(1) Andrea Loparic already obtained in 1977, announced in [23] and finally published in [24], a semantical characterization of Johannson-Kolmogorov minimal logic and intuitionistic propositional logic $\mathbf{IPL}$ in terms of a $2$-valued, non-truth-functional valuation semantics which induces a decision procedure for these logics. Similar to Grätz’s method (and to the adaptation to $\mathbf{IPL}$ proposed here), Loparic’s approach is based on a criterion for distinguishing between what she called semivaluations and valuations, the latter being the useful ones. This would correspond to level valuations, since such valuations are associated with $\varphi$-saturated sets, and they are the ones that characterize the logic. The associated algorithm generates a non-deterministic truth-table for a given formula, as in the cases of S4 and $\mathbf{IPL}$ presented in this paper. Then, using formal criteria based on implicative formulas, it deletes certain rows. Thus, a formula is a tautology of the given logic if and only if it receives the value 1 in all remaining rows.
(2) J. Kearns proposed in [18] a semantics for (first-order) intuitionistic logic based on the notion of justifications. A 3-valued Nmatrix, whose domain is not formed by truth-values but by values of justification (++ for ‘justified’, $o$ for ‘weakly unjustified’, and - for ‘strongly unjustifiable’), is considered for interpreting the connectives of $\mathbf{IPL}$. A sentence is weakly unjustified if it is not justified by the currently justified sentences but might become justified later. In turn, strongly unjustifiable sentences are those already ruled out by the currently justified sentences. Given two valuations $v$ and $v^{\prime}$ over the Nmatrix, $v\preceq v^{\prime}$ ($v$ is contained in $v^{\prime}$) if $v^{\prime}$ preserves the values ++ (and so, the values -) assigned by $v$ to each formula. Hence, if $v^{\prime}(\alpha)=o$ then $v(\alpha)=o$, for any $\alpha$. The intuitive idea is that $v^{\prime}$ is a ‘possible future’ of $v$ (hence $v^{\prime}$ has more ‘knowledge’ than $v$, provided that the weak epistemic status $o$ may be corrected to the stronger ++ or - in the future). Kearns constructed a hierarchy of level valuations (as he would later do for modal logics) starting with level 0, which is formed by all valuations on the Nmatrix. The $(n+1)$th-level valuations are the $n$th-level valuations satisfying certain restrictions defined in terms of the sets $X_{n}^{v}=\{v^{\prime}\mid v^{\prime}$ is an $n$th-level valuation such that $v\preceq v^{\prime}\}$, for any $n$th-level valuation $v$. A basic feature is that, if $v$ is an $(n+1)$th-level valuation and $v^{\prime}(\alpha)\neq+$ for every $v^{\prime}\in X_{n}^{v}$ then $v(\alpha)=$-. That is, if $\alpha$ is not justified in any possible future of $v$, then $\alpha$ must be strongly unjustifiable at present $v$. In a second paper [20] on the subject, published two years after [18] and seven months after [19], Kearns improved the interpretation of the justification semantics and simplified its formal presentation by using trees instead of levels. The contained relation $\preceq$ is crucial to define the trees whose nodes are valuations. It should be stressed that Kearns’s proposal constitutes a 3-valued RNmatrix semantics for $\mathbf{IPL}$, although it differs significantly from our approach. In particular, Kearns’s RNmatrix does not seem to induce a decision procedure for $\mathbf{IPL}$.
(3) Concerning modal logics, in [21], right after the publication of [12], O. Lahav and Y. Zohar proposed a decision procedure for systems K and KT based on 4- and 3-valued RNmatrices, respectively, by using a decision procedure different (but closely related) to Grätz’s approach. Indeed, for ${\bf L}\in\{{\bf K},{\bf KT}\}$ they considered a sequent calculus $\mathcal{G}_{\bf L}$ and sets $\mathbb{V}_{\bf L}^{\mathcal{F},n}$ of $n$th-level valuations relative to a set $\mathcal{F}$ of formulas (the domain of such valuations) and $n$ is the maximum number of applications of the necessitation rule validated by such valuations. Given a sequent $\Gamma\Rightarrow\varphi$, the expression $\vdash_{\mathcal{G}_{\bf L}}^{\mathcal{F},n}\Gamma\Rightarrow\varphi$ denotes that there exists a derivation in $\mathcal{G}_{\bf L}$ of $\Gamma\Rightarrow\varphi$ with a derivation containing exclusively formulas in $\mathcal{F}$ and such that the number of applications of the inference rule of necessitation $\Delta\Rightarrow\psi/\Box\Delta\Rightarrow\Box\psi$ in any of the branches of the derivation is limited by $n$. They proved that $\vdash_{\mathcal{G}_{\bf L}}^{\mathcal{F},n}\Gamma\Rightarrow\varphi$ if and only if $\varphi$ is a consequence of $\Gamma$ w.r.t. $n$th-level valuations in $\mathbb{V}_{\bf L}^{\mathcal{F},n}$. Moreover, they obtained a decision procedure for K by considering, for each sequent $\Gamma\Rightarrow\varphi$ to be tested, the set of level valuations $\mathbb{V}_{\bf L}^{\mathcal{F},n}$, where $\mathcal{F}$ is the set of subformulas of $\Gamma\cup\{\varphi\}$ and $n=4^{|\mathcal{F}|}$. For KT, it is enough considering $n=3^{|\mathcal{F}|}$, since the RNmatrix for KT is 3-valued. Moreover, that RNmatrix coincides with the one proposed in [12] for KT. It is worth noting that [21] considers the reduced version of the underlying Nmatrix, expanded to the full signature. This produces the Nmatrix $\mathcal{M}^{\prime}_{S4}$ for S4 of Table 2, but where $\Box^{\prime}(2)=\{1,2\}$. Despite these similarities, the decision procedure in [21] (as mentioned earlier) differs from that in [12]. It consists of generating the set of $n$th-level valuations over $\mathcal{F}$ for appropriate $n$ and $\mathcal{F}$.
(4) More recently, A. Solares-Rojas proposed a 3-valued Nmatrix semantics for $\mathbf{IPL}$ based on intuitionistic Kripke models and the novel notion of depth-bounded approximations (see [29, Chapter 5]). Within this semantic approach to $\mathbf{IPL}$, a 3-valued Nmatrix — whose truth-values are intuitionistic truth (1), falsity (0), and indeterminacy ($u$) — gives the meaning of the connectives in terms of information that a given agent holds. The valuations over the Nmatrix aim to model actual information states, ordered by a certain refinement relation $\sqsubseteq_{3}$. This relation models how the agent’s knowledge evolves when new information from reliable external sources is acquired. The semantics for $\mathbf{IPL}$ is defined via intuitionistic Kripke frames expanded with a family of valuations indexed by the worlds, where the refinement order corresponds to the frame’s accessibility relation. Some similarities between this approach and Kearns’s (mentioned in point (2)) may be observed. Conceptually, both approaches aim to model the evolution of states of knowledge: the information state of an agent in [29], and the knowledge of an idealized community in [18, 20]. Regarding technical aspects, by identifying $1$, 0, and $u$ with ++, -, and $o$ respectively, the Nmatrix in [29, Chapter 5] coincides with that in [18], with one exception: while $o\vee o=\{o,+\}$ in Kearns’s work, $u\vee u=\{u\}$ (i.e., $o\vee o=\{o\}$) in Solares-Rojas’s approach. Additionally, the refinement relation $\sqsubseteq_{3}$ between valuations coincides with the contained relation $\preceq$ introduced in [18]. The relationship between both approaches (in particular, the connections with the improved version [20], which is closer to Kripke and Beth’s semantics) deserves to be investigated. Regarding the connections with our approach, although the 3-valued Nmatrices proposed in [29, Chapter 5] and the one we gave in Definition 8.4 are quite different, it is not immediate whether the semantics proposed by Solares-Rojas could be seen as given by an RNmatrix. This is the subject of future investigation, as well as the relationship between our approach and the semantics of justification introduced by Kearns.
12 Concluding remarks
In this paper, we proposed a new decision procedure for intuitionistic propositional logic ($\mathbf{IPL}$) based on a 3-valued restricted Nmatrix semantics. This procedure was obtained by abstracting the composition of Grätz’s decision procedure for S4 [12] and Gödel’s faithful translation from $\mathbf{IPL}$ to S4. A 3-valued Nmatrix with an intuitive interpretation in epistemic terms was proposed for $\mathbf{IPL}$. By adapting Grätz’s algorithm for S4, a suitable notion of level valuations was obtained, which allowed us to define a straightforward algorithm for deleting spurious rows in the non-deterministic truth-tables generated by the Nmatrix for any formula. The soundness and completeness proofs for the RNmatrix and the resulting truth-table algorithm for $\mathbf{IPL}$ were presented in full technical detail.
It is worth noting that the Nmatrix for $\mathbf{IPL}$ introduced in Definition 8.4 has a clear and intuitive interpretation in terms of knowledge (see comments after Definition 8.4) and constructibility. Indeed, recall that the valuation associated with a $\varphi$-saturated set $\Delta$ in $\mathbf{IPL}$ (given in Definition 8.14) acts as a 3-valued characteristic function of $\Delta$, and note the observations in Remark 8.15.
By considering finite and decidable RNmatrix semantics as a generalization of standard finite-valued deterministic truth-tables, the results presented here provide a way to overcome the limitations obtained by Gödel and Dugundji for $\mathbf{IPL}$ and S4, respectively, as mentioned in Section 1.
As a topic of future research, we intend to explore the meaning of the partial valuations corresponding to the rows of the non-deterministic truth-tables in $\mathbf{IPL}$ and S4. In particular, we are studying the possibility of generating finite Kripke models for $\mathbf{IPL}$, KT and S4 from the obtained truth-tables. This could be connected with finite approximability ([5, p. 49 and Ch. 11]) or with mini-canonical models ([13, Ch. 8]). Finally, the extension of this methodology to other modal logics, as well as Prawitz’s ecumenical systems (see [27]), will also be investigated. Some steps in this direction were taken in [22].
Acknowledgments: The first and second authors thank the support of the São Paulo Research Foundation (FAPESP, Brazil) through the PhD scholarship grant #21/01025-3 and the Thematic Project RatioLog #20/16353-3. The second author also acknowledges support from the National Council for Scientific and Technological Development (CNPq, Brazil), through the individual research grant #309830/2023-0. The third author thanks Fundação Carlos Chagas de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ, Brazil), through the research grant #E-26.210.971/2019.